OpenStudy (anonymous):

Dont understand steps on how to do this problem. 2y-4x=24 x=3y - 1

OpenStudy (anonymous):

You have to do the same steps you followed in your last question

sam (.sam.):

2y-4x=24.......(1) x=3y - 1........(2) Substitute (2) into (1) 2y-4x=24 2y-4(3y - 1)=24 y=-2 From (2), x=3y - 1 x=3(-2)-1 x=-7

OpenStudy (anonymous):

you already know x = 3y - 1 so substitute it into the other eqution and solve for y

OpenStudy (anonymous):

Do what @.Sam. says:)

OpenStudy (anonymous):

@jazy and @.Sam. okay thanks, im a little rusty with this lol been awhile.

OpenStudy (anonymous):

It all right it took me a while too:) After a time it'll be easier to understand...personal experience lol:)

OpenStudy (anonymous):

perfect ! thanks sam.